Exercise 3.9:

Constants A physics book slides off a horizontal table fop with a speed of 1.70 m/s. It strikes the floor after a time of 0 450 s Ignore air resistance.

media%2Fe93%2Fe9362787-8b69-4675-93d7-c1

media%2Ffac%2Ffac05d10-47b5-496e-bad2-a9

- Part E Find the magnitude of the books velocity just before the book reaches the floor Express your answer with the appropriate units. Units Submit Request Answer Part F Find the direction of the book s velocity just before the book reaches the floor below the horicontal
Answer #1

(a) y = (1/2)gt^2

= (1/2) (9.8 m/s^2) (0.450 s)^2 = 0.992 m

(b) x = vt = (1.70 m/s) (0.450 s) = 0.765 m

(c) still 1.70 m/s

(d) v = v0 + at = 0 + (-9.8 m/s^2)(0.450 s) = -4.41 m/s

(e) sqrt(1.7^2 + 4.41^2) = 4.73 m/s

(f) arctan (-4.41/1.7) = -68.9 degrees, i.e., 68.9 degrees below horizontal